1. What is integral of tan x dx?

2. What is d/dx sin x?

3. What is integral of e^3x dx?

4. What is d/dx e^7x?


1
Expert's answer
2021-04-27T09:16:28-0400
  1. "\int" tan x dx = - ln |Cos x| + C = ln |sec x| + C.
  2. "\dfrac{dsinx}{dx}=cosx ."
  3. "∫e^{3x} dx=e^{3x} \dfrac{1}{3} + c"
  4. 7"e^{7x}" .